Find the vertex of the graph of the quadratic function. Determine whether the graph opens upward or​ downward, find any​ intercepts, and sketch the graph.
f(x)=-x^2-2x+3
A. The vertex is:
B. The parabola opens downward or upward
C. The x-intercepts are:
D. The Y-intercepts are:
E; Graph

Answer by stanbon(75887)

Find the vertex of the graph of the quadratic function. Determine whether the graph opens upward or​ downward, find any​ intercepts, and sketch the graph.
f(x)=-x^2-2x+3
A. The vertex is:
x = -b/(2a) = 2/(-2) = -1
y = f(-1) = -1+2+3 = 4
------------------------------------------
B. The parabola opens downward or upward
Down because a = -1
---------------------------
C. The x-intercepts are:
Solve -x^2-2x+3 = 0
x^2 + 2x - 3 = 0
(x+3)(x-1) = 0
x = -3 or x = 1
Intercepts:: (-3,0) ; (1,0)
----------------------------------
D. The Y-intercepts are:
f(0) = -0^2-2*0+3 = 3
Intercept:: (0,3)
---------------------------
E; Graph
Plot the points and sketch the parabola.
-----------
Cheers,
Stan H.
-----------------